OpenStudy (anonymous):

Can someone point me in the right direction on how to solve this proof? Note: I'm in PreCalc and these are written as sqr roots not exponents. (6^1/2 - 2^1/2)/4 = [(2 - 3^1/2)^1/2]/2

OpenStudy (anonymous):

let x=(6^1/2-2^1/2)/4 or x=2^1/2(3^1/2-1)/4 Squaring x^2=2(3+1-2*3^1/2)/16 x^2=(4-2*3^1/2)/8 or x^2=2(2-3^1/2)/8 or x^2=(2-3^1/2)/4 or x={(2-3^1/2)^1/2}/2 =R.H.S